# One Wave to Explain Them All: A Unifying Perspective on Post-hoc Explainability

### v1

@@ B.2.2 Results
Table 4 reports the Insertion and Deletion (Petsiuk et al., 2018) scores compared to other attribution methods. We can see that the good results reported in Table 1 are mostly driven by our method’s very good results in Insertion. We can see that WAM systematically outperforms the other methods in both Insertion and Deletion.
Table 4: Insertion and Deletion (Petsiuk et al., 2018) scores obtained on 1,000 images from the validation set of ImageNet and for different model architectures. For insertion, higher is better, and for deletion, lower is better. Bolded results are the best and underlined values are the second best.
Method ResNet ConvNext EfficientNet DeiT Mean
@@
Audio.
Table 5 displays the results of the quantitative evaluation of WAM when the metrics are computed from the mel-spectrogram. We can see that we still achieve state-of-the-art performance for the Faithfulness of spectra and are in line with other metrics for the Input Fidelity. These results highlight the added value of computing the metrics from the wavelet coefficients rather than from the mel-spectrogram.
Table 5: Evaluation scores of WAM and comparison with baselines on 400 audios from the first fold of ESC-50. The column ”ESC” indicates that the samples are unaltered. The column ”+WN” indicates that the samples have 0 dB Gaussian white noise. We report the results with explanations generated from the mel-spectrogram of the waveform. Bolded results are the best and underlined values are the second best.
Method Faithfulness ($\uparrow$) Insertion ($\uparrow$) Deletion($\downarrow$) FF ($\uparrow$) Fid-In ($\uparrow$)
ESC50 +WN ESC50 +WN ESC50 +WN ESC50 +WN ESC50 +WN
IntegratedGradients 0.264 0.310 0.267 0.312 0.047 0.045 0.207 0.207 0.220 0.225
GradCAM 0.072 0.073 0.274 0.274 0.201 0.201 0.137 0.135 0.517 0.542
Saliency 0.066 0.065 0.220 0.221 0.154 0.156 0.166 0.168 0.253 0.245
WAMSG (ours) 0.009 0.007 0.169 0.166 0.159 0.161 0.152 0.149 0.117 0.122
WAMIG (ours) 0.000 0.004 0.168 0.171 0.168 0.167 0.149 0.149 0.105 0.128
B.3 Randomization check
Movitation.
The sanity checks introduced by Adebayo et al. (2018) aim at assessing whether an explanation depends on the model’s parameters and the input labels. These tests aim to assess the faithfulness of an explanation beyond visual evaluation. The randomization test evaluates whether an explanation depends on the model’s parameters. Parameters have a strong effect on a model’s performance. Therefore, for a saliency method to be useful for debugging or analyzing a model, it should be sensitive to its parameters. Adebayo et al. (2018) proposed different methods to randomize the model parameters. One particularly interesting implementation is the “cascading” randomization, in which the weights are randomized from the top to the bottom layers.
Method and results.
We compute WAM for our 1,000 ImageNet validation samples for a set of increasingly randomized models. A randomized layer is a layer that we reset at its initial value. We consider a ResNet-18 and randomize its layers from the shallowest conv1 to the deepest fc. We then compute the rank correlation (or Pearson correlation coefficient, Pearson, 1896) between the WAM of the original, fully-trained model (labeled orig) and the randomized models (labeled by the name of the layer until which they are randomized).
Figure 8 presents the results. The dotted line represents the average rank correlation across the 1,000 images, and the intervals represent the 95% confidence intervals. We can see that the correlation between WAM s significantly decreases as the randomization increases, thereby showing that WAM is sensitive to the model’s parameters. The lower decrease that we observe for WAMIG compared to WAMSG comes from the fact that WAMIG reflects more the inter-scales distribution of the importance than WAMSG does. As pointed out by Rahaman et al. (2019) and Yin et al. (2019), even random models exhibit a spectral bias, i.e., a natural tendency to favor lower frequencies over higher ones, which translates here into the fact of naturally putting more importance on coarser scales rather than finer scales, no matter the depth of the randomization.
Illustrations.
Figure 9, Figure 21, Figure 22, Figure 23 display qualitative examples to further assess how WAM varies when the randomization depth of the model increases. We can see that the explanation is no longer informative when the model is randomized.
Appendix C Additional results
C.1 Frequency-centric perspectives on model robustness
Scales in the wavelet domain correspond to dyadic frequency ranges in the Fourier domain. Several works documented a correlation between the reliance on low frequency to make predictions and the robustness of the model (Zhang et al., 2022; Chen et al., 2022; Wang et al., 2020). We can leverage WAM to characterize a model’s robustness, thus connecting feature attribution and robustness. On Figure 10, we evaluate the reliance on the different scales by summing the importance of each component within each scale. We average this importance over 1,000 images and thus obtain the average importance of each scale for a model’s prediction. We compare a vanilla ResNet-50 (He et al., 2016) with three adversarially robust models : ADV (Madry et al., 2018), ADV-Fast (Wong et al., 2020) and ADV-Free (Shafahi et al., 2019). We can see that adversarially robust models rely more on the coarsest scale (leftmost bars on Figure 10) than the vanilla ResNet-50. On the other hand, they rely less on the finest scales (i.e., the highest frequencies, corresponding to the rightmost bars on Figure 10), thus backing the existing results established in the Fourier domain. Therefore, WAM can be used to characterize the robustness of a model.
Figure 10: Assessment of model’s robustness with WAM.
C.2 Additional results on the minimal images
Complements on “optimal” minimal images.
@@ C.2 Additional results on the minimal images
Figure 11 illustrates the effect of the parameter $\alpha$ on the sparsity of the minimal images. We can see that the stronger $\alpha$, the sparser the image, but at the expense of a higher logit value. We can see that the first components of the image that disappear are the background, then the colors and eventually the shape of the target class.
@@
Static minimal images.
Another method used to derive minimal images is to directly sort the gradient coefficients obtained with WAM and reconstruct the original image using decreasingly important coefficients.
We referred to the “insertion” quantile the threshold value after which the prediction of the model was correct, meaning that the partially reconstructed image contained enough information. We also considered the original image for completeness and removed gradually important coefficients. It turned out that both quantiles coincided, thus highlighting the fact that WAM identifies the necessary and sufficient amount of information for a correct prediction.
Figure 12 displays an example of a sufficient image extracted using WAM. We can see that the background information is unnecessary, contrary to details around the fox’s ear and eyes. Reconstruction artifacts are caused by the fact that we independently ranked the coefficients across the three color channels.
Figure 12: Extraction of the necessary and sufficient information for a correct prediction with WAM. In the example, we can see that only 22% of the components are necessary, the remaining being redundant.
Figure 13 presents illustrations of minimal images obtained using WAM. We can see that the background can be filtered out, while foreground details are essential. The minimal images have been obtained using information solely from WAM.
Figure 13: Example of minimal images (rightmost column), which contain enough information to correctly predict the image label. Compared to the input image (leftmost column), the background can be disregarded, but high-resolution details are necessary on the animals. The column in the center plots the wavelet heatmap obtained with WAM.
C.3 Overlap experiment on audio classification
Figure 14 illustrates that WAM is able to filter relevant parts corrupted or mixed audio signals. In addition, it highlights the key part of the target signal without requiring any training. Figure 14 qualitatively illustrates application of WAM for audio signals. Herein, we perform an overlap experiment to mix a corrupting audio with a target audio to form the input audio. The model’s prediction does not alter after introducing the corruption, and thus, the model is expected to still rely on parts of input audio coming from the target audio for its decision. The interpretation audio in Figure 14 generated using top wavelet coefficients provides insights into the decision process and supports this hypothesis. In particular, it almost entirely filters out corruption audio, and without requiring any training, it also clearly emphasizes key parts of target audio.
Figure 14: Qualitative illustration of WAM for audio via an Overlap experiment. The audio of the target class (’Crow’) is mixed with a corrupting audio (’Chirping birds’) to form the input to the classifier. Interpretation audio reconstructed with important wavelet coefficients virtually eliminate signal from the corrupting audio, and also clearly emphasize parts of the target class audio (indicated with green boxes).
C.4 Additional plots
Additional illustrations on images.
Figure 15 presents additional illustrations of WAM for images. We illustrate the important coefficients (using the IntegratedGradients variant for smoothing the gradients) in the wavelet domain (second column) as a heatmap on the original image (third column), and we show the important components on the image, i.e., what the model needs to see based on the WAM on the fourth column.
Figure 15: Additional visualizations of WAM for images. First column: original image. Second column: WAM in the wavelet domain. Third column: heatmap in the pixel domain. Fourth column: Filtered image illustrating the image regions that need to be well defined for the model to predict the image label correctly.
Additional minimal images.
Figure 16 presents additional examples of minimal images. We can see that the color information does not appear as important for maximizing the model’s prediction. On the other hand, the texture and edge information are essential. It would be interesting to replicate this method on a shape-biased model such as those proposed by Chen et al. (2020); Geirhos et al. (2019) to see whether the behavior remains the same or not.
@@ C.4 Additional plots
Additional explanations on shapes.
Figure 17, Figure 18, Figure 19 and Figure 20 present additional visualization on shapes.
Figure 17: Decomposition of the different important scales on a voxel with WAM.
Figure 18: Decomposition of the different important scales on a voxel with WAM.
Figure 19: Decomposition of the different important scales on a voxel with WAM.
Figure 20: Heatmaps combining the importance at different scales on different voxels from 3D MNIST.
Additional randomization checks.
Figure 21, Figure 22, and Figure 23 present additional examples of randomization checks.
Figure 21: Illustration of the randomization test as a cascade of randomizations of the layers of the classifier. From left to right shows the explanation from the WAM for an increasingly randomized ResNet-18.
Figure 22: Illustration of the randomization test as a cascade of randomizations of the layers of the classifier. From left to right shows the explanation from the WAM for an increasingly randomized ResNet-18.
Figure 23: Illustration of the randomization test as a cascade of randomizations of the layers of the classifier. From left to right shows the explanation from the WAM for an increasingly randomized ResNet-18.
Additional audio overlap experiments.
Figure 24 and Figure 25 present additional examples of overlap experiments on waveforms.
@@
Figure 24: Qualitative illustration of WAM for audio via an Overlap experiment. The audio of the target class (’Dog’) is mixed with a corrupting audio (’Cat’) to form the input to the classifier. Interpretation audio reconstructed with important wavelet coefficients virtually eliminates signal from the corrupting audio and clearly emphasizes parts of the target class audio (indicated with green boxes).
Figure 25: Qualitative illustration of WAM for audio via an Overlap experiment. The audio of the target class (’Dog’) is mixed with a corrupting audio (’Rooster’) to form the input to the classifier. Interpretation audio reconstructed with important wavelet coefficients virtually eliminates signal from the corrupting audio and clearly emphasizes parts of the target class audio (indicated with green boxes).

### v2

@@ C.3 Additional evaluation results
Additional methods.
Table 5 compares WAM with more recent feature attribution methods, considering a ResNet as a model backbone. We compare WAM against LayerCAM (Jiang et al., 2021), Guided Backpropagation (Springenberg et al., 2014), LRP (Binder et al., 2016) and SRD (Han et al., 2024).
Table 5: Evaluation of WAM on ImageNet with recent feature attribution methods using a ResNet backbone. Best results are bolded, second best underlined.
Metric Faithfulness ($\uparrow$) Insertion ($\uparrow$) Deletion ($\downarrow$) $\mu$-Fidelity ($\uparrow$)
Guided Backprop 0.001 0.090 0.089 0.216
LayerCAM 0.139 0.432 0.293 0.143
LRP 0.039 0.277 0.238 0.126
SRD 0.002 0.102 0.101 1.151
WAMIG (ours) 0.438 0.557 0.119 0.215
WAMSG (ours) 0.344 0.344 0.078 0.170
Volumes
Complementary results on MNIST3D.
In Table 6 we evaluate our method against Saliency (Zeiler & Fergus, 2014a), GradCAM (Selvaraju et al., 2017), GradCAM++ (Chattopadhay et al., 2018), IntegratedGradients (Sundararajan et al., 2017), GuidedBackpropagation (Springenberg et al., 2014) and SmoothGrad (Smilkov et al., 2017) on the dataset MNIST3D. We report the Insertion, the Deletion, and the Faithfulness. Overall, WAM performs as well as or better than the other methods.
Table 6: Evaluation scores (3D) on the full test set of MNIST3D. Best results are bolded, second best underlined.
Metric Faithfulness ($\uparrow$) Insertion ($\uparrow$) Deletion($\downarrow$)
Saliency -0.166 0.108 0.274
GradCAM -0.225 0.106 0.330
GradCAM++ -0.250 0.104 0.354
IntegratedGradients -0.123 0.108 0.232
GuidedBackprop -0.110 0.108 0.218
SmoothGrad -0.190 0.107 0.297
WAMSG (ours, $J=2$) -0.153 0.108 0.261
WAMIG (ours, $J=2$) -0.094 0.106 0.200
WAMSG (ours, $J=1$) -0.152 0.108 0.260
WAMIG (ours, $J=1$) -0.096 0.107 0.203
Evaluation on MedMNIST.
MedMNIST (Yang et al., 2023) is a collection of standardized biomedical 2D and 3D images. We evaluated WAM on two datasets from MedMNIST, AdrenalMNIST and VesselMNIST. Table 7 shows the evaluation results of WAM on AdrenalMNIST and VesselMNIST using different model topologies.
Table 7: Evaluation scores (3D) on the full test set of AdrenalMNIST and VesselMNIST. Best results are bolded, second best underlined.
Model ResNet3D EfficientNet3D 3D Former
Metric Faith ($\uparrow$) Ins ($\uparrow$) Del ($\downarrow$) Faith ($\uparrow$) Ins ($\uparrow$) Deletion ($\downarrow$) Faith ($\uparrow$) Ins ($\uparrow$) Del ($\downarrow$)
AdrenalMNIST Saliency -0.136 0.633 0.769 -0.085 0.679 0.764 -0.027 0.715 0.742
GradCAM -0.195 0.529 0.724 -0.052 0.693 0.745 -0.065 0.679 0.744
GradCAM++ -0.137 0.531 0.668 -0.076 0.695 0.771 -0.091 0.656 0.748
IntegratedGradients -0.204 0.531 0.735 -0.056 0.705 0.761 -0.077 0.666 0.743
Guided Backpropagation -0.132 0.585 0.717 0.003 0.683 0.680 -0.004 0.685 0.689
SmoothGrad 0.115 0.663 0.548 -0.070 0.686 0.756 -0.051 0.680 0.731
WAMSG (ours, $J=2$) -0.279 0.467 0.746 0.049 0.666 0.617 0.070 0.718 0.648
WAMIG (ours, $J=2$) -0.224 0.407 0.631 0.086 0.684 0.598 0.106 0.719 0.613
WAMSG (ours, $J=1$) -0.273 0.474 0.746 0.026 0.661 0.635 0.050 0.717 0.667
WAMIG (ours, $J=1$) -0.224 0.478 0.702 0.013 0.699 0.686 0.026 0.697 0.671
VesselMNIST Saliency -0.068 0.833 0.901 -0.747 0.131 0.878 -0.099 0.762 0.861
GradCAM -0.095 0.789 0.884 -0.728 0.126 0.854 -0.065 0.804 0.869
GradCAM++ 0.011 0.871 0.860 -0.725 0.124 0.849 -0.066 0.797 0.863
IntegratedGradients -0.116 0.771 0.887 -0.743 0.127 0.870 -0.060 0.808 0.868
GuidedBackprop -0.229 0.654 0.883 -0.717 0.121 0.838 0.002 0.843 0.841
SmoothGrad 0.137 0.861 0.724 -0.674 0.130 0.804 0.045 0.851 0.806
WAMSG (ours, $J=2$) -0.054 0.848 0.902 -0.674 0.126 0.800 0.020 0.798 0.778
WAMIG (ours, $J=2$) -0.064 0.756 0.820 -0.630 0.126 0.756 0.010 0.793 0.783
WAMSG (ours, $J=1$) -0.070 0.832 0.902 -0.692 0.128 0.820 0.001 0.799 0.798
WAMIG (ours, $J=1$) -0.069 0.787 0.856 -0.691 0.128 0.819 -0.070 0.764 0.834
Appendix D Additional use cases
D.1 Revisiting meaningful perturbations
Meaningful perturbations in the wavelet domain.
Meaningful perturbations (Fong & Vedaldi, 2017; Fong et al., 2019) identify the most relevant regions of an input for a model’s prediction by learning an optimal mask that hides as little of the image as possible while significantly altering the model’s output. This is done by optimizing a mask through gradient-based updates to minimize the classifier’s confidence while maintaining a smoothness constraint. The result highlights the most important features in a way that is more structured and localized than saliency maps. We revisit this framework by proposing to recast the problem in the wavelet domain as a more suitable space for optimization. The wavelet domain inherently captures spatial and spectral information, providing a natural structure for producing meaningful and interpretable solutions. Specifically, we solve the following optimization problem:
$\bm{m}^{\star}=\operatorname*{arg\,min}_{\bm{m}\in[0,1]^{|\mathcal{X}|}}\bm{f}_{c}\left(\mathcal{W}^{-1}(\bm{z}\odot\bm{m})\right)+\alpha\|\bm{m}\|_{1},$ (19)
$\bm{m}_{i+1}=\bm{m}_{i}-\eta\nabla_{\bm{m}_{i}}\left(\bm{f}_{c}\left(\mathcal{W}^{-1}(\bm{z}\odot\bm{m}_{i})\right)+\alpha\|\bm{m}_{i}\|_{1}\right),$ (20)
Figure 11: A) Sparsity-optimized minimal images. We revisit meaningful perturbation by optimizing the sparsity of the wavelet transform using masking, instead of optimizing the mask in pixel space. The displayed examples show that the resulting minimal images reveal the model’s reliance on textures. B) Sparsity Pareto front. As $\alpha$ increases, the sparsity of the wavelet coefficients increases (x-axis), but beyond a certain point, too much information is lost and the logit score drops to zero. However, we observe that many components can be removed before adversely affecting the model. Results are averaged across 1,000 images optimized for 500 steps and for $\alpha$ ranging in $[0,100]$ for each image.
Traditional meaningful perturbation methods (Fong & Vedaldi, 2017) focus on spatial localization, identifying clusters of pixels that answer the question of where the important features are located. However, this spatial emphasis alone provides a limited understanding of the underlying data structure. In contrast, by operating in the wavelet domain, our method captures both the what – the relevant scales – and the where – their spatial locations. This dual information enriches the explanation by revealing the location and the nature of the features influencing the model’s decision. These results also show that we qualitatively recover the results from Kolek et al. (2023).
Effect of the regularizer on the sparsity of the images.
@@ D.1 Revisiting meaningful perturbations
Figure 12 illustrates the effect of the parameter $\alpha$ on the sparsity of the minimal images. We can see that the stronger $\alpha$, the sparser the image, but at the expense of a higher logit value. We can see that the first components of the image that disappear are the background, then the colors, and eventually the shape of the target class.
@@
Minimal images and applications.
Figure 13 presents additional examples of minimal images. We can see that the color information does not appear as important for maximizing the model’s prediction. On the other hand, the texture and edge information are essential. It would be interesting to replicate this method on a shape-biased model such as those proposed by Chen et al. (2020) or Geirhos et al. (2019) to see whether the behavior remains the same or not.
@@
D.2 Multi-class classification
Evaluation in a multi-class setting.
We evaluate WAM in a multi-class setting to assert whether our method is still informative when an image depicts different objects. To evaluate our method, we rely on the Pointing Game (Zhang et al., 2018) benchmark. The Pointing Game assesses the spatial precision of saliency or attribution maps. For each image, the location of the maximum activation in the saliency map (the “point”) is compared against the ground-truth object location. If the point lies within the ground-truth bounding box (or segmentation mask), it is considered a “hit”; otherwise, it is a “miss”.
The final accuracy is computed as:
Pointing Game $\text{Pointing Game Accuracy}=\frac{\#\text{Hits}}{\#\text{Hits}+\#\text{Misses}}.$
This metric focuses on whether the most confident localization prediction corresponds to the true object location, without requiring full object delineation. We evaluate WAM alongside competing methods (GradCAM, GradCAM++, and SmoothGrad) on the PASCAL VOC 2012 test set (Everingham et al., 2010), a widely used benchmark for visual recognition tasks. The dataset contains 20 object classes and 1456 test images. Each image is annotated with object bounding boxes and class labels, making it suitable for evaluating weakly-supervised localization via the Pointing Game.
Table 8: Pointing game experiment results. Evaluation carried out on the test set of the PASCAL VOC 2012 test split. Best results are bolded and second best underlined.
Model ResNet EfficientNet ConvNext VGG
GradCAM 0.476 0.454 0.528 0.437
GradCAM++ 0.476 0.454 0.528 0.437
SmoothGrad 0.429 0.432 0.430 0.430
WAMSG (ours) 0.606 0.584 0.584 0.538
WAMIG (ours) 0.604 0.601 0.632 0.586
Table 8 presents the results: we can see that WAM outperforms the competing baselines at the Pointing Game, thereby showing its ability to identify the relevant spatial parts of the input image for a given class. This further backs the fact that the wavelet domain is informative in multi-category images and suggests that WAM  can be adapted to object detectors.
Visualizations.
Once shown that WAM can effectively distinguish objects in multi-category settings, we can analyze which regions are highlighted as important by WAM on such images. In Figure 14, we observe that the WAM successfully highlights “cat-related coefficients,” as the cat’s nose and its corresponding regions in the wavelet domain are more active in the rightmost image. Interestingly, although less prominent, parts of the dog’s face also remain significant. This suggests that while the spatial separation between the cat and dog is clear in the original image, the distinction becomes less apparent in the wavelet domain.
Figure 14: Illustration of multi-category images. Example on an image depicting a cat and a dog to highlight that the WAM points towards the regions corresponding to the target label (Dogs on the right, Cats on the left).
D.3 Overlap experiment in audio classification
Figure 15 illustrates that WAM is able to filter relevant parts corrupted or mixed audio signals. In addition, it highlights the key part of the target signal without requiring any training. Figure 15 qualitatively illustrates application of WAM for audio signals. Herein, we perform an overlap experiment to mix a corrupting audio with a target audio to form the input audio. The model’s prediction does not alter after introducing the corruption, and thus, the model is expected to still rely on parts of input audio coming from the target audio for its decision. The interpretation audio in Figure 15 generated using top wavelet coefficients provides insights into the decision process and supports this hypothesis. In particular, it almost entirely filters out the corruption audio, and without requiring any training, it also clearly emphasizes key parts of target audio.
@@ D.3 Overlap experiment in audio classification
Figure 15: Qualitative illustration of WAM for audio via an Overlap experiment. The audio of the target class (“Crow”) is mixed with a corrupting audio (“Chirping birds”) to form the input to the classifier. Interpretation audio reconstructed with important wavelet coefficients virtually eliminates signal from the corrupting audio, and also clearly emphasizes parts of the target class audio (indicated with green boxes).
D.4 Scales meet semantics: the case of remote sensing applications
The indexation in scales of the wavelet coefficients finds a natural use case in remote sensing applications (Kasmi et al., 2023b), where scales correspond to actual physical objects, as the pixels correspond to distances on the ground. Therefore, given an aerial image whose ground sampling distance is 20 cm/pixel, the 1-2 pixel scale (the finest details), corresponds to details that have a size comprised between 20 to 40 cm.
On Figure 16, we consider a case where a classification model is trained on a source domain and deployed on a target domain, thus mimicking situations where a model is used for regular updates. We can see that the OOD image is slightly different from the source image, in the sense that it is less noisy. While one would expect the model’s prediction not to change, it turns out that the photovoltaic (PV) panel is no longer recognized on the rightmost image. Attribution in the pixel domain only is not very informative as to why this happens. Turning into the wavelet domain, we can better grasp why the model no longer detects the PV panel. The background noise, located mostly in the finest scales, is no longer present on the OOD image, and thus the PV panel is no longer recognized. Kasmi et al. (2025) discussed how wavelet-based attribution methods help improve the reliability of classifiers deployed in such operational settings.
Figure 16: Application case of WAM to remote sensing. While both images depict the same photovoltaic panel, a model no longer detects the panel when evaluated on images coming from a different distribution (”OOD” or ”out-of-distribution” image, compared to images from the source distribution). Attribution in the wavelet domain thanks to the WAM gives hints as to why the PV system is no longer recognized: on the image of the original domain, which is more noisy, the model relies on noise at the finest scales that are no longer present on the OOD image. Thus suggesting that the representation learned by the model lacks robustness. Based on Kasmi et al. (2023b)